A pendulum of length 0.99m is taken to the moon by an astronaut. The period of the pendulum is 4.9s. what is the value of g on the surface of the moon?


1
Expert's answer
2022-01-06T09:06:53-0500

The value of g is


"\dfrac{4\pi^2 l}{T^2}=1.63\text{ m/s}^2."